OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is parallel to the line shown y = -2x + 2 y = 2x - 14 y = 1/2x - 8 y = -1/2x - 4

OpenStudy (anonymous):

Need it urgent!

OpenStudy (anonymous):

Which point and which line?

OpenStudy (anonymous):

that one

OpenStudy (anonymous):

Really Need Help

OpenStudy (anonymous):

y = 1/2x-8

OpenStudy (anonymous):

Oo wait..it is y=-1/2x +4

OpenStudy (anonymous):

That is not an option

OpenStudy (anonymous):

The slope for the line given in the diagram = (-5)/10 = -1/2 in the options only one line has that slope i.e option d

OpenStudy (anonymous):

y=-1/2x - 4 My typing is bad today!

OpenStudy (anonymous):

THANK YOU VERY MUCH!

OpenStudy (anonymous):

Can you help me on this? too?

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6.

OpenStudy (anonymous):

y = 1/3x - 8 y = -3x + 12 y = 3x - 24 y = -1/3x - 4